OpenStudy (141823):

A rectangle has a width that is 1/8 its length. If the perimeter of the rectangle is 126 inches, what is the width of the rectangle?

OpenStudy (141823):

@ranga

Directrix (directrix):

Let the length of the rectangle be x. Then, the width is (1/8)x. The perimeter of the rectangle is two times the length plus two times the width. p = (1/8) x + (1/8)x + x + x 126 = (1/8) x + (1/8)x + x + x Solve the equation ^^ above for x Multiply x by (1/8) to get the width. @141823

OpenStudy (141823):

I'm not sure if this is correct, but I got x= 123.75 so 123.75*.25=30.9375

OpenStudy (141823):

@Directrix

Directrix (directrix):

I got something different for x. @141823 What equation did you get after you combined the like terms?

Directrix (directrix):

What did you get for this: (1/8) x + (1/8)x + x + x

OpenStudy (141823):

Wait, I think I did it wrong, let me try again

OpenStudy (141823):

(.125)x + (.125)x + x + x = 126, I'm not sure what to do from now on

OpenStudy (141823):

@Directrix

OpenStudy (141823):

@ranga help

OpenStudy (141823):

Nevemind i got it

Directrix (directrix):

(.125)x + (.125)x + x + x = 126 2.25x = 126 x = ? @141823 Multiply x by (1/8) to get the width.